OpenStudy (anonymous):

Help with functions?? I am having trouble solving this because I am not good with fractions! f(1/3)= 4(1/3)^2-5(1/3)+2

OpenStudy (anonymous):

[4*(1/3)*(1/3)] - [5*(1/3)] + 2 = (4*1/9) - (5/3) + 2 (from here you find the least common denominator, which is 9. therefore, every fraction will be over 9 4/9 is already over 9 5/3-> 15/9, as you multiply by 3/3 (which is the same as multiplying by 1) 2 -> 18/9, as you multiply by 9/9 (which is the same as multiplying by 1) = 4/9 - 15/9 + 18/9 = 7/9

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

any time!